### v3

# Quasi-one-dimensional spin transport in altermagnetic $Z^{3}$ nodal net metals

Abstract
In three dimensions, quasi-one-dimensional (Q1D) transport has traditionally been associated with systems featuring a Q1D chain structure. Here, based on first-principle calculations, we go beyond the common belief to show that the Q1D transport can also be realized in many three-dimensional (3D) altermagnetic (AM) metals with a topological nodal net in momentum space but lacking Q1D chain structure in real space, including the existing compounds $\beta$-Fe2(PO4)O, Co2(PO4)O, and LiTi2O4. These materials exhibit an AM ground state and feature an ideal crossed $Z^{3}$ Weyl nodal line in each spin channel, formed by three straight and flat nodal lines traversing the entire Brillouin zone. These nodal lines eventually lead to an AM $Z^{3}$ nodal net. Surprisingly, longitudinal conductivity $\sigma_{xx}$ in these topological nodal net metals is dozens of times larger than $\sigma_{yy}$ and $\sigma_{zz}$ in the up-spin channel, while $\sigma_{yy}$ dominates transport in the down-spin channel. This suggests a distinctive Q1D transport signature in each spin channel, with orthogonal principal moving directions for the two spin channels, resulting in Q1D direction-dependent spin transport. This novel phenomenon cannot be found in both conventional 3D bulk materials and Q1D chain materials. In particular, it gradually disappears as the Fermi level moves away from the nodal net, further confirming its topological origin. Our work not only enhances the comprehension of topological physics in altermagnets but also opens a new direction for the exploration of topological spintronics.
The coupling of magnetism and topological states has sparked extensive research interest [1, 2, 3, 4, 5]. Magnetic topological materials exhibit a plethora of novel physical phenomena, including the quantum anomalous Hall effect [6, 7, 8], anomalous Hall effect [9, 10, 11, 12], anomalous Nernst effect [13, 14, 15, 16, 12], and magnetoresistance effect [17, 18], which hold promising application prospects in spintronics. Magnetic ordering, on one hand, diminishes the system’s symmetry by breaking time-reversal symmetry (${\cal T}$) and certain crystal symmetries, presenting challenges in extending topological phases to magnetic systems [19, 20, 21, 22]. On the other hand, it introduces an additional spin degree of freedom, paving the way for achieving electric control of spin, a critical pursuit in spintronics [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
In recent years, significant progress has been made in realizing various topological phases within ferromagnetic (FM) systems. Notable examples include the Co3Sn2S2 [37, 13], Heusler compound Co2MnGa [14, 38], and rutile-type metal fluorides LiV2F6 [39], rhombohedral transition metal trifluorides PdF3 [12], etc. These compounds exhibit distinct physical properties within the magnetic space group, such as Weyl points and nodal lines, along with anomalous magnetic transport phenomena like the giant anomalous Hall effect and anomalous Nernst effect. Moreover, the emerging field of topological altermagnetic (AM) spintronics, facilitated by AM topological states, has garnered considerable attention [25, 26, 28, 29, 27]. Unlike FM systems, AM systems lack a macro-magnetic moment, offering advantages such as eliminating stray magnetic fields, reducing power requirements, and improving operating frequency response [40, 41, 42]. Notably, various topological phases have been achieved in both AM materials and conventional antiferromagnets [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53], paving the way for novel applications in spintronics.
A unique form of topological nodal line, termed the $Z^{3}$ nodal line, has been proposed in nonmagnetic systems [54]. Unlike conventional nodal lines, this structure traverses the entire Brillouin zone (BZ) and is characterized by three integers $Z^{3}=(n_{x},n_{y},n_{z})$, indicating the number of times the line winds around each corresponding direction [54, 55]. Remarkably, owing to significant anisotropy, the ideal $Z^{3}$ nodal line is another structure besides layered and chain structures that can lead to low-dimensional transport phenomena. Here, the “ideal” means the $Z^{3}$ nodal line is straight and has a small energy variation, as illustrated in Fig. 1. For instance, an ideal $Z^{3}$ nodal line with $n_{z}=1$ would result in quasi-two-dimensional (Q2D) transport, with slower electron velocity along the $z$-direction compared to the $x$-$y$ plane [see Fig. 1(a)]. With certain symmetries, two $Z^{3}$ nodal lines may form a crossed $Z^{3}$ nodal line, as shown in Fig. 1(b), where the $y$-component longitudinal conductivity will dominates the transport when the Fermi level is around the nodal line. This means that the crossed $Z^{3}$ nodal line also can lead to Q1D transport signatures. However, unlike layered and chain structures defined in real space, the $Z^{3}$ nodal line is defined in momentum space. Therefore, a natural question then is: Does an ideal $Z^{3}$ nodal line or crossed $Z^{3}$ nodal line exist in AM materials? If so, what novel transport phenomena distinguish them from conventional nodal line semimetals and metals with layered or chain structures.
Figure 1: Illustration of (a) an ideal $Z^{3}$ nodal line with $n_{z}=1$, and (b) an ideal crossed $Z^{3}$ nodal line with $n_{z}=1$ and $n_{x}=1$. The figures from left to right denote the shape, the band structure and the Fermi surface of $Z^{3}$ and $Z^{3}$ crossed nodal lines. Since the Fermi surface of $Z^{3}$ and $Z^{3}$ crossed nodal lines respectively feature cylindrical and flattened shapes, the two nodal lines should exhibit Q2D and Q1D transport properties.
In this work, we demonstrate that AM materials, such as $\beta$-Fe2(PO4)O, Co2(PO4)O, and LiTi2O4, are ideal topological $Z^{3}$ nodal net metals. Significantly, these materials deviate from conventional layered or Q1D chain materials. In these compounds, each spin channel exhibits a crossed $Z^{3}$ nodal line, formed by a $Z^{3}$ nodal line in the $k_{z}=0$ plane and two $Z^{3}$ nodal lines with $n_{z}=1$. At the BZ boundary, the two spin-polarized crossed nodal lines contact each other, resulting in a novel AM $Z^{3}$ nodal net configuration, as depicted in Fig. 3(c). Previous studies have elucidated the electronic structures of $\beta$-Fe2(PO4)O and LiTi2O4 [56, 57]. Additionally, $\beta$-Fe2(PO4)O has been identified as an X-type altermagnetic material [58], characterized by sublattice-selective spin-polarized transport and spin torque. However, the impact of the AM $Z^{3}$ nodal net on the spin transport properties of these systems remains unexplored.
Based on the first-principles calculations, we have discovered a class of materials characterized by novel Q1D spin transport properties. These materials exhibit an ideal $Z^{3}$ nodal net near the Fermi level. Furthermore, their longitudinal conductivity exhibits Q1D features, where the conductivity $\sigma_{xx}$ ($\sigma_{yy}$) is approximately two orders of magnitude larger than $\sigma_{yy}$ ($\sigma_{xx}$) and $\sigma_{zz}$ in the up-spin (down-spin) channel, leading to novel Q1D direction-dependent spin transport. Notably, away from the AM $Z^{3}$ nodal net, the spin conductivity remains anisotropic but gradually loses its Q1D features, suggesting a close relationship between this novel spin transport and the AM $Z^{3}$ nodal net. To illustrate our findings, we provide concrete examples and explore their transport properties in depth.
Example 1: $\beta$-Fe2(PO4)O and Co2(PO4)O.–The material $\beta$-Fe2(PO4)O is a mixed-valence iron oxyphosphate featuring Fe2+/Fe3+ oxidation states [59, 60, 61]. Its single crystal was synthesized in the 1980s and experimentally confirmed to exhibit an AM phase with a Néel temperature as high as 408 K [61].
[…]
Figure 2: (a) The conventional cell and (b) primitive cell of the $\beta$-Fe2(PO4)O compound with AM ordering. Red and blue arrows represent up-spin and down-spin magnetic moments, respectively. (c) The bulk Brillouin zone (BZ) with the high symmetry points labeled.
Since the spin-orbit coupling (SOC) in $\beta$-Fe2(PO4)O is negligible [62], we consider the spin-resolved band structure of the material without SOC, as illustrated in Fig. 3(a). In the $k_{z}=0$ plane, $\beta$-Fe2(PO4)O features a $Z^{3}$ Weyl nodal line in each spin channel, protected by the glide mirror symmetries $G_{z}=\{M_{z}|\frac{1}{2}0\frac{1}{2}\}$. The profiles of these two Weyl nodal lines are depicted in Fig. 3(b), showing that the Weyl nodal line in the up-spin (down-spin) channel is very straight and traverses the BZ in the $k_{y}$ ($k_{x}$) direction. The two spin-resolved Weyl nodal lines contact with each other at the $X$ point in the BZ. Furthermore, for both spin channels, the two bands around the Fermi level are degenerate along two inequivalent $P$-$X$ paths, resulting in another two $Z^{3}$ nodal lines with $n_{z}=1$, which are not only straight but also flat. These $Z^{3}$ nodal lines ultimately form an AM $Z^{3}$ nodal net, as depicted in Fig. 3(c).
Figure 3: (a) Electronic band structure and the projected density of states (PDOS) of $\beta$-Fe2(PO4)O. The energy region where the $Z^{3}$ AM nodal net appears is highlighted by green. (b) The shapes of $Z^{3}$ nodal line in each spin channel. (c) The shape of $Z^{3}$ nodal net in the BZ. Red and blue lines represent up-spin and down-spin Weyl nodal lines.
As all the $Z^{3}$ Weyl nodal lines in $\beta$-Fe2(PO4)O are ideal and exhibit an extremely strong anisotropic feature, one can expect that the longitudinal conductivity $\sigma_{xx}^{\uparrow}$ ($\sigma_{yy}^{\downarrow}$) of the up-spin (down-spin) channel in this material should be much larger than $\sigma_{yy}^{\uparrow}$ ($\sigma_{xx}^{\downarrow}$) and $\sigma_{zz}^{\uparrow}$ ($\sigma_{zz}^{\downarrow}$). In the absence of SOC, spin serves as a good quantum number, allowing for the conductivity of up-spin and down-spin electrons to be accurately defined and calculated. Therefore, to investigate the unique Q1D direction-dependent spin transport properties of the AM $Z^{3}$ nodal net in $\beta$-Fe2(PO4)O, we calculate the spin-resolved longitudinal conductivities for different directions, and the obtained results are shown in Fig. 4. Three key features are observed.
Firstly, the conductivities for each spin are not independent but are connected by $C_{4z}\mathcal{T}$ symmetry, leading to $\sigma_{xx}^{\uparrow}=\sigma_{yy}^{\downarrow}$, $\sigma_{yy}^{\uparrow}=\sigma_{xx}^{\downarrow}$, and $\sigma_{zz}^{\uparrow}=\sigma_{zz}^{\downarrow}$.
Secondly, around the AM nodal net, the system’s conductivity exhibits a novel spin-dependent signature where the longitudinal conductivity $\sigma_{xx}^{\uparrow}$ ($\sigma_{yy}^{\downarrow}$) is dozens of times larger than the other two longitudinal conductivities. This implies that while the electric current is isotropic, the spin current is strongly anisotropic and shows a Q1D signature in each spin channel. This phenomenon is distinct from systems with a chain structure, where both electric and spin currents would exhibit a Q1D signature. Since the dominant transport directions for up-spin and down-spin are orthogonal, $\beta$-Fe2(PO4)O can be utilized to generate a current with strong spin polarization, and the spin polarization can be easily switched by rotating the sample 90 degrees along the $z$ axis.
Thirdly, away from the $Z^{3}$ AM nodal net, the spin conductivity remains anisotropic but progressively loses the Q1D signature. For high-energy electrons, the $\sigma_{xx}$ ($\sigma_{yy}$) of the up-spin (down-spin) channel and the other two conductivities are of the same order of magnitude (see Fig. 4). This strongly suggests that the spin-polarized Q1D transport is closely associated with the $Z^{3}$ nodal net rather than the system’s structure.
Figure 4: The calculated conductivity of the $\beta$-Fe2(PO4)O compound in (a) up-spin channel and (b) down-spin channel. The $Z^{3}$ AM nodal net appears in the green shaded region.
The synthesis of Co2(PO4)O has also been reported [63], showing a similar crystalline structure, magnetic configuration, and electronic band properties as $\beta$-Fe2(PO4)O [62]. Notably, the Q1D spin transport signature is also significant in Co2(PO4)O. Further details can be found in Ref. [62].
Example 2: LiTi2O4.–In addition to the $\beta$-Fe2(PO4)O compound, the LiTi2O4 compound also exhibits an AM $Z^{3}$ nodal net near the Fermi level. The lattice structure of LiTi2O4, depicted in Fig. 5(a), belongs to a cubic Bravais lattice, which differs from that of $\beta$-Fe2(PO4)O. The AM ground state of LiTi2O4 is characterized by a magnetic symmetry of 141.554 ($I4^{\prime}_{1}$/$am^{\prime}d$) [57].
The SOC effect in LiTi2O4 is also negligible. Figure 5(c) illustrates the spin-resolved band structure of LiTi2O4 in the absence of SOC. Scanning through the Brillouin zone (BZ) reveals the profile of the AM $Z^{3}$ nodal net, as depicted in Fig. 5(b). This AM nodal net configuration consists of one nodal line in the $k_{z}=0$ plane and two nodal lines along two inequivalent $P$-$X$ paths for each spin channel. The obtained spin longitudinal conductivities are displayed in Fig. 5(d-e), exhibiting similar features to those observed in $\beta$-Fe2(PO4)O and Co2(PO4)O. This further confirms that Q1D spin transport is a common characteristic of materials possessing an ideal $Z^{3}$ nodal net.
Figure 5: (a) Crystal structures of the LiTi2O4 compound. Red and blue arrows represent up-spin and down-spin magnetic moments, respectively. (b) BZ and the shape of the $Z^{3}$ nodal net. Red and blue lines represent up-spin and down-spin Weyl nodal lines. (c) Electronic band structure and the PDOS of LiTi2O4 compound. The calculated conductivity in (d) up-spin channel and (e) down-spin channel. The $Z^{3}$ AM nodal net appears in the green shaded region.
Lattice model of AM $Z^{3}$ nodal net.– To further understand the Q1D spin transport, we establish a simple effective lattice model for the AM $Z^{3}$ nodal net, based on $\beta$-Fe2(PO4)O. We consider a 3D tetragonal body-centred lattice with an AM ordering, and assume that this lattice has the same MSG (No. 141.554) as the $\beta$-Fe2(PO4)O. The unit cell of the antiferromagnet includes an even number of lattice sites: up spins on half of sites and down spins on the remaining half. The lattice has four atoms with $s$-like orbitals per unit cell at $8d$ Wyckoff sites (labeled as $\{A,B,C,D\}$), which correspond to the positions occupied by the Fe atoms of $\beta$-Fe2(PO4)O [see Fig. 2(b)].
The lattice has space group (SG) symmetry $D_{4h}^{19}$ (e.g. the SG 141), which is generated by a four-fold screw-rotation $S_{4z}=\{C_{4z}|\frac{1}{4}\frac{-1}{4}\frac{1}{4}\}$, a two-fold rotation ${C}_{2x}$ and a spatial inversion ${P}$. Combining $S_{4z}^{2}$ and ${P}$ leads to a glide mirror plane $G_{z}=\{M_{z}|\frac{1}{2}0\frac{1}{2}\}$. Without altermagnetic ordering, the lattice also has time-reversal symmetry ${\cal T}$. In the absence of SOC, using the four sites as a basis, the symmetry operators can be represented as
$S_{4z}$ $=\left[\begin{array}[]{cccc}0&0&1&0\\ 0&0&0&1\\ 0&1&0&0\\ 1&0&0&0\end{array}\right],C_{2x}=\left[\begin{array}[]{cccc}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{array}\right],$ (9)
$P=\Gamma_{0,0}$ and ${\cal T}=\Gamma_{0,0}{\cal K}$ with $\Gamma_{i,j}=\sigma_{i}\otimes\sigma_{j}$. Here, $\sigma_{0}$ is identity matrix and $\bm{\sigma}$ denote the Pauli matrixes. The corresponding model Hamiltonian under these symmetry constraints may be written as [64, 65]
${\cal H}_{0}$ $=\left[\begin{array}[]{cc}H_{1}&H^{\prime}\\ H^{\prime\dagger}&H_{2}\end{array}\right],$ (12)
where $H_{1}=t_{1}\cos\frac{k_{x}+k_{z}}{2}\sigma_{1}$, $H_{2}=t_{1}\cos\frac{k_{y}+k_{z}}{2}\sigma_{1}$ and
$H^{\prime}$ $=t_{2}\left[\begin{array}[]{cc}\cos\frac{k_{z}}{2}&\cos\frac{k_{y}}{2}\\ \cos\frac{k_{x}}{2}&\cos\frac{k_{x}+k_{y}+k_{z}}{2}\end{array}\right],$ (15)
with $t_{i}$ ($i=1,2$) the real parameters.
Then, we introduce spin degree of freedom and turn on a AM ordering along $z$-direction [see Fig. 2(a)]. The AM ordering breaks ${\cal T}$, $S_{4z}$ and $C_{2x}$, while holds $G_{z}$, $S_{2z}$, $P$ and the combined operation $S_{4z}{\cal T}$ and $C_{2x}{\cal T}$. To the leading order, the AM Hamiltonian may be written as
${\cal H}$ $=$ ${\cal H}_{0}s_{0}+J_{z}s_{z}\Gamma_{3,0},$ (16)
with $\bm{s}$ the Pauli matrixes acting on spin space and $J_{z}$ denoting the strength of the AM potential. When the AM potential ($J_{z}>0$) is much stronger than other hopping and energy terms, the eight bands are divided into two groups with each group including four bands. Since the two groups are separated in energy space by $2J_{z}$, the model (16) can be simplified as two effectively four-band model [66]. The basis of the upper subsystem is $(|A\uparrow\rangle,|B\uparrow\rangle,|C\downarrow\rangle,|D\downarrow\rangle)^{T}$ and that for the lower subsystem is $(|A\downarrow\rangle,|B\downarrow\rangle,|C\uparrow\rangle,|D\uparrow\rangle)^{T}$. We focus on the upper subsystem and omit $J_{z}$ from it, as it only shifts the bands. Following the classical method [67], the upper four-band model can be expressed as
${\cal H}_{\rm{eff}}$ $=\frac{t^{2}_{2}}{4J_{z}}+\left[\begin{array}[]{cc}H_{1}&0\\ 0&H_{2}\end{array}\right]+\frac{t^{2}_{2}}{4J_{z}}\left[\begin{array}[]{cc}H_{1}^{\prime}&0\\ 0&H_{2}^{\prime}\end{array}\right],$ (21)
with
$H_{1}^{\prime}({\bm{k}})$ $=2(\cos\frac{k_{x}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{x}+k_{y}+k_{z}}{2})\sigma_{1}$
$+\left[\begin{array}[]{cc}\cos k_{y}+\cos k_{z}&0\\ 0&\cos k_{x}+\cos(k_{x}+k_{y}+k_{z})\end{array}\right],$ (24)
and $H_{2}^{\prime}(k_{x},k_{y},k_{z})=H_{1}^{\prime}(k_{y},k_{x},k_{z})$.
The band structure of ${\cal H}_{\rm{eff}}$ is calculated and plotted in Ref. [62], which qualitatively reproduces the band structure of $\beta$-Fe2(PO4)O. Our simple four-band effective model hence can serve as a good starting point for the further study of the physics of AM $Z^{3}$ nodal net.
Conclusions.–In conclusion, we have demonstrated the feasibility of achieving the ideal AM $Z^{3}$ nodal net in real materials and shown that the novel Q1D spin transport phenomena inherent to this topological structure. We have detailed the mechanisms whereby electrons with up-spin predominantly flow in one direction, while those with down-spin move primarily in a perpendicular direction, a characteristic distinctly contrasting with conventional 3D bulk materials and Q1D chain materials. This behavior could be effectively detected through the measurement of anisotropic spin resistance, offering a practical approach to exploring the unique physics of AM systems. Our work not only provides a solid foundation for exploring the novel physics of AM systems but also holds great promise for applications in topological spintronics.