### v4

# Quasi-One-Dimensional Spin Transport in Altermagnetic $Z^{3}$ Nodal Net Metals

Abstract
In three dimensions, quasi-one-dimensional (Q1D) transport has traditionally been associated with systems featuring a Q1D chain structure. Here, based on first-principle calculations, we go beyond this understanding to show that the Q1D transport can also be realized in certain three-dimensional (3D) altermagnetic (AM) metals with a topological nodal net in momentum space but lacking Q1D chain structure in real space, including the existing compounds $\beta$-Fe2(PO4)O, Co2(PO4)O, and LiTi2O4. These materials exhibit an AM ground state and feature an ideal crossed $Z^{3}$ Weyl nodal line in each spin channel around Fermi level, formed by three straight and flat nodal lines traversing the entire Brillouin zone. These nodal lines eventually lead to an AM $Z^{3}$ nodal net. Surprisingly, the electronic conductivity $\sigma_{xx}$ in these topological nodal net metals is dozens of times larger than $\sigma_{yy}$ and $\sigma_{zz}$ in the up-spin channel, while $\sigma_{yy}$ dominates transport in the down-spin channel. This suggests a distinctive Q1D transport signature in each spin channel, and the principal moving directions for the two spin channels are orthogonal, resulting in Q1D direction-dependent spin transport. This novel phenomenon cannot be found in both conventional 3D bulk materials and Q1D chain materials. In particular, the Q1D spin transport gradually disappears as the Fermi energy moves away from the nodal net, further confirming its topological origin. Our work not only enhances the comprehension of topological physics in altermagnets but also opens a new direction for the exploration of topological spintronics.
The coupling of magnetism and topological states has sparked extensive research interest [1, 2, 3, 4, 5]. Magnetic topological materials exhibit a plethora of novel physical phenomena, including the quantum anomalous Hall effect [6, 7, 8], anomalous Hall effect [9, 10, 11, 12], anomalous Nernst effect [13, 14, 15, 16, 12], and magnetoresistance effect [17, 18], which hold promising application prospects in spintronics. Magnetic ordering, on one hand, diminishes the system’s symmetry by breaking time-reversal symmetry (${\cal T}$) and certain crystal symmetries, presenting challenges in extending topological phases to magnetic systems [19, 20, 21, 22]. On the other hand, it introduces an additional spin degree of freedom, paving the way for achieving electric control of spin, a critical pursuit in spintronics [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33].
In recent years, significant progress has been made in realizing various topological phases within ferromagnetic and antiferromagnetic systems. Notable examples include PdF3 [12], Co3Sn2S2 [34, 13], Co2MnGa [14, 35], and LiV2F6 [36], Mn3Sn/Ge [37], CaCrO3 [38], YbMnBi2 [16], etc. These magnetic compounds exhibit interesting topological states, such as Weyl points and nodal lines, along with anomalous transport phenomena like the giant anomalous Hall effect and anomalous Nernst effect.
Remarkably, a new class of magnetic order: altermagnetism has been proposed very recently [39, 40, 41, 42, 43], and has garnered considerable attention. The altermagnetic (AM) systems feature collinear-compensated magnetic order, and alternating spin polarizations in both crystal structure and electronic band structure [42]. Particularly, due to large spin splitting and robustness against magnetic field perturbations, the AM materials have demonstrated intriguing spin-dependent transport phenomena, leading to AM spintronics [40, 41, 44]. Some topological phases have been achieved in AM materials (they were recognized as antiferromagnetic materials based on the previous classification of magnetic phases) [45, 46, 47], offering new material choices and operational mechanisms for the design and application of spintronic devices.
A unique form of topological nodal line, termed the $Z^{3}$ nodal line, has been proposed in nonmagnetic systems [48]. Unlike conventional nodal lines, this structure traverses the entire Brillouin zone (BZ) and is characterized by three integers $Z^{3}=(n_{x},n_{y},n_{z})$, indicating the number of times the line winds around each corresponding direction [48, 49]. Remarkably, owing to significant anisotropy, the ideal $Z^{3}$ nodal line is another structure besides layered and chain structures that can lead to low-dimensional transport phenomena. Here, the “ideal” means the $Z^{3}$ nodal line is straight and has a small energy variation, as illustrated in Fig. 1. For instance, an ideal $Z^{3}$ nodal line with $n_{z}=1$ would result in quasi-two-dimensional (Q2D) transport, with slower electron velocity along the $z$-direction compared to the $x$-$y$ plane [see Fig. 1(a)]. With certain symmetries, two $Z^{3}$ nodal lines may form a crossed $Z^{3}$ nodal line, as shown in Fig. 1(b), where the $y$-component longitudinal electronic conductivity will dominates the transport of system, suggesting a Q1D transport in topological crossed $Z^{3}$ nodal line semimetals. However, unlike layered and chain structures defined in real space, the $Z^{3}$ and crossed $Z^{3}$ nodal line are defined in momentum space. Particularly, due to the altermagnetism, the crossed $Z^{3}$ nodal lines in AM materials generally come in pair with different $Z^{3}$ index, forming an AM $Z^{3}$ nodal net. This is a unique property that cannot be found in ferromagnetic and antiferromagnetic materials. Therefore, a natural question then is: Does an ideal $Z^{3}$ nodal line or crossed $Z^{3}$ nodal line exist in AM materials? If so, what novel transport phenomena distinguish them from the materials with layered or chain structures.
Figure 1: Illustration of (a) an ideal $Z^{3}$ nodal line with $n_{z}=1$, and (b) an ideal crossed $Z^{3}$ nodal line with $n_{z}=1$ and $n_{x}=1$. The figures from left to right denote the shape, the band structure and the Fermi surface of the nodal lines. Since the Fermi surface of $Z^{3}$ and $Z^{3}$ crossed nodal lines respectively feature cylindrical and flattened shapes, the two nodal lines should exhibit Q2D and Q1D transport properties.
In this work, we find that the AM materials: $\beta$-Fe2(PO4)O, Co2(PO4)O, and LiTi2O4, are ideal topological $Z^{3}$ nodal net metals, featuring two crossed $Z^{3}$ nodal lines around Fermi level in both spin channels [see Fig. 3(c)]. Previous studies have elucidated the electronic structures of $\beta$-Fe2(PO4)O and LiTi2O4 [50, 51]. Additionally, $\beta$-Fe2(PO4)O has been identified as a X-type altermagnetic material [52]. However, the impact of the AM $Z^{3}$ nodal net on the transport properties of these systems remains unexplored.
Based on the first-principles calculations, we demonstrate that the ideal AM $Z^{3}$ nodal net indeed can generate an unique transport phenomenon, namely, in the up-spin and down-spin channels, the electronic conductivity around Fermi level are dominated by $\sigma_{xx}^{\uparrow}$ and $\sigma_{yy}^{\downarrow}$, respectively, leading to novel Q1D direction-dependent spin transport. Notably, away from the AM $Z^{3}$ nodal net, the spin conductivity remains anisotropic but gradually loses its Q1D features, suggesting a close relationship between this novel spin transport and the AM $Z^{3}$ nodal net.
Figure 2: (a) The conventional and (b) primitive cell of $\beta$-Fe2(PO4)O. Red and blue arrows represent up-spin and down-spin magnetic moments, respectively. (c) denotes the BZ.
Example 1: $\beta$-Fe2(PO4)O and Co2(PO4)O.–The material $\beta$-Fe2(PO4)O is a mixed-valence iron oxyphosphate featuring Fe2+/Fe3+ oxidation states [53, 54, 55]. It has been experimentally identified as an antiferromagnetic material with a Néel temperature as high as 408 K [55] in the 1980s, but according to the modern classification for magnetic systems [42], it is an AM material [52].
[…]
Figure 3: (a) Electronic bands and projected density of states (PDOS) of $\beta$-Fe2(PO4)O. The energy region where the $Z^{3}$ AM nodal net appears is highlighted by green. (b) The shapes of the $Z^{3}$ crossed nodal lines in up-spin and down-spin channels. The two $Z^{3}$ crossed nodal lines form a $Z^{3}$ nodal net. (c) The shape of the $Z^{3}$ nodal line on the $k_{z}$ = 0 plane for each spin channel indicated by the arrows in (b). The color map indicates the local gap between the two bands that construct the $Z^{3}$ nodal line.
Since the spin-orbit coupling (SOC) in $\beta$-Fe2(PO4)O is negligible [57], we consider the spin-resolved band structure of the material without SOC, which is plotted in Fig. 3(a). Interestingly, in each spin channel, there exist two bands around the Fermi level that are degenerate along two inequivalent $P$-$X$ paths, resulting in two ideal $Z^{3}$ nodal lines with $n_{z}=1$, as the lines are not only straight but also flat [see Fig. 3(a,b)]. Moreover, in the $k_{z}=0$ plane, $\beta$-Fe2(PO4)O features a $Z^{3}$ Weyl nodal line in both spin channels, protected by the glide mirror symmetries ~$\tilde{{\cal M}}_{z}$. The profiles of the Weyl nodal lines are numerically calculated and depicted in Fig. 3(c), showing that the Weyl nodal lines in both spin channels are very straight and traverse the BZ. Thus, the up-spin (down-spin) channel exhibits a $Z^{3}$ crossed nodal line with $n_{z}=1$ and $n_{y}=1$ ( $n_{x}=1$), as depicted in Fig. 3(b). Furthermore, the two spin-resolved $Z^{3}$ crossed nodal lines contact with each other at the $X$ point in the BZ, ultimately forming an AM $Z^{3}$ nodal net. A detailed symmetry analysis of such a $Z^{3}$ nodal net is provided in the Supplemental Material (SM) [57].
As all the $Z^{3}$ Weyl nodal lines in $\beta$-Fe2(PO4)O are ideal, one can expect that the longitudinal conductivity $\sigma_{xx}^{\uparrow}$ ($\sigma_{yy}^{\downarrow}$) of the up-spin (down-spin) channel in this material should be much larger than $\sigma_{yy}^{\uparrow}$ ($\sigma_{xx}^{\downarrow}$) and $\sigma_{zz}^{\uparrow}$ ($\sigma_{zz}^{\downarrow}$). In the absence of SOC, the electronic conductivity of up-spin and down-spin electrons can be accurately and independently calculated. Therefore, to investigate the unique Q1D direction-dependent spin transport properties of the AM $Z^{3}$ nodal net in $\beta$-Fe2(PO4)O, we calculate the spin-resolved longitudinal electronic conductivities for both spin channels. The obtained results are shown in Fig. 4, from which three key features are observed.
Figure 4: The calculated electronic conductivity (left axis) and $\eta$ (right axis) of the $\beta$-Fe2(PO4)O compound in (a) up-spin and (b) down-spin channels. The $Z^{3}$ AM nodal net appears in the green shaded region. Here, we set temperature of system as $T=50$ K. Hence, the electronic conductivity is finite in the band gap.
Firstly, the electronic conductivity in each spin generally is anisotropic that $\sigma_{xx}^{\uparrow(\downarrow)}\neq\sigma_{yy}^{\uparrow(\downarrow)}\neq\sigma_{zz}^{\uparrow(\downarrow)}$. However, the electronic conductivities for the two spin channels are not independent, but have $\sigma_{xx}^{\uparrow}=\sigma_{yy}^{\downarrow}$, $\sigma_{yy}^{\uparrow}=\sigma_{xx}^{\downarrow}$, and $\sigma_{zz}^{\uparrow}=\sigma_{zz}^{\downarrow}$, consistent with the symmetry analysis.
Secondly, we find that the anisotropy of the electronic conductivity in each spin channel is extremely enhanced around the AM nodal net (Fermi level), where $\sigma_{xx}^{\uparrow}$ ($\sigma_{yy}^{\downarrow}$) is dozens of times larger than the other two longitudinal conductivities. To quantitatively describe it, we define a dimensionless quantity
$\eta=\frac{\sigma_{xx}^{\uparrow}-\sigma_{\perp}^{\uparrow}}{\sqrt{(\sigma_{xx}^{\uparrow})^{2}+(\sigma_{\perp}^{\uparrow})^{2}}}=\frac{\sigma_{yy}^{\downarrow}-\sigma_{\perp}^{\downarrow}}{\sqrt{(\sigma_{yy}^{\downarrow})^{2}+(\sigma_{\perp}^{\downarrow})^{2}}},$ (1)
with $\sigma_{\perp}^{\uparrow(\downarrow)}={\sqrt{(\sigma_{yy(xx)}^{\uparrow(\downarrow)})^{2}+(\sigma_{zz}^{\uparrow(\downarrow)})^{2}}}$. $\eta=1$ corresponds to the strongest anisotropy and ideal Q1D direction-dependent spin transport, where $\sigma_{xx}^{\uparrow}$ and $\sigma_{yy}^{\downarrow}$ completely dominate the electronic transport in up-spin and down-spin channels, respectively. As shown in Fig. 4, $\eta$ approaches 1 only in the energy range where the AM nodal net appears. We have further checked that in this energy range, the Fermi surface of the system indeed has a flattened shape [57].
Remarkably, the Q1D transport in $\beta$-Fe2(PO4)O is distinct from systems with chain structures. For the former, although the spin current has Q1D characteristic, the electric current is isotropic in the $x$-$y$ plane, as $\sigma_{xx}^{\uparrow}+\sigma_{xx}^{\downarrow}=\sigma_{yy}^{\uparrow}+\sigma_{yy}^{\downarrow}$. In contrast, for the latter, both electric and spin currents exhibit Q1D signature. Moreover, due to the direction-dependent Q1D spin transport, $\beta$-Fe2(PO4)O can be utilized to generate a current with strong spin polarization, and the spin polarization can be easily switched by rotating the sample 90 along $z$ axis.
Thirdly, the PDOS in Fig. 3(a) shows that the energy bands from -1 eV to 3 eV are mainly contributed by the Fe atoms. However, away from the $Z^{3}$ AM nodal net, the spin conductivity loses the Q1D signature, as $\eta$ rapidly deviates from 1 in the lower energy ($<-0.8$ eV), and also is far from 1 in the higher energy ($>1.8$ eV), as shown in Fig. 4. This strongly suggests that the spin-polarized Q1D transport is closely associated with the $Z^{3}$ nodal net rather than the structural arrangement of the atoms in $\beta$-Fe2(PO4)O.
The synthesis of Co2(PO4)O has also been reported [58], showing a similar crystalline structure, magnetic configuration, electronic bands and the Q1D spin transport as $\beta$-Fe2(PO4)O [57].
Example 2: LiTi2O4.–The LiTi2O4 also exhibits an AM $Z^{3}$ nodal net near the Fermi level. The lattice structure of LiTi2O4 without magnetic momentum belongs to a cubic Bravais lattice, which differs from that of $\beta$-Fe2(PO4)O. However, the ground state of LiTi2O4 has AM configuration, as depicted in Fig. 5(a). Since the Néel vector is incompatible with the crystal symmetry, the MSG of the LiTi2O4 is calculated as No. 141.554 [56, 51]. The optimized lattice parameters of the conventional cell are $a=b=6.053$ and $c=8.753$.
Figure 5(c) shows the electronic bands of the LiTi2O4 without SOC, as SOC effect here is negligible. Scanning through the whole BZ reveals the existence of an ideal AM $Z^{3}$ nodal net [see Fig. 5(b)], which share similar configuration as that in $\beta$-Fe2(PO4)O. Particularly, the direction-dependent Q1D spin transport is observed in LiTi2O4, and again only appears in the energy where AM $Z^{3}$ nodal net appears [see Fig. 5(c-e)]. This further confirms that Q1D spin transport is a common characteristic of materials possessing an ideal $Z^{3}$ nodal net.
Figure 5: (a) Crystal structures of LiTi2O4. Red and blue arrows represent up-spin and down-spin magnetic moments, respectively. (b) BZ and the shape of the $Z^{3}$ crossed nodal line in each spin channel. (c) Electronic bands and PDOS of LiTi2O4. (d,e) The calculated electronic conductivity (left axis) and $\eta$ (right axis) in (d) up-spin and (e) down-spin channels. The $Z^{3}$ AM nodal net appears in the green shaded region.
Lattice model of AM $Z^{3}$ nodal net.– To further understand the Q1D spin transport, we establish a simple effective lattice model for the AM $Z^{3}$ nodal net, based on $\beta$-Fe2(PO4)O. We consider a 3D tetragonal body-centred lattice with an AM ordering, and assume that this lattice has the same MSG (No. 141.554) as the $\beta$-Fe2(PO4)O. The unit cell of the system includes an even number of lattice sites: up spins on half of sites and down spins on the remaining half. The lattice has four atoms with $d_{x^{2}-y^{2}}$-like orbitals [57] per unit cell at $8d$ Wyckoff sites (labeled as $\{A,B,C,D\}$), which correspond to the positions occupied by the Fe atoms of $\beta$-Fe2(PO4)O [see Fig. 2(b)].
The lattice belongs to SG No. 141, which is generated by a four-fold screw-rotation $\tilde{{\cal C}}_{4z}$ = $\{C_{4z}|\frac{1}{4}\frac{-1}{4}\frac{1}{4}\}$, a two-fold rotation ${C}_{2x}$, and spatial inversion ${\cal P}$. Without altermagnetic ordering, the lattice also has time-reversal symmetry ${\cal T}$. In the absence of SOC, using the four sites as a basis, the symmetry operators can be represented as $\tilde{{\cal C}}_{4z}=(\Gamma_{+,0}+\Gamma_{-,1})/2$, $C_{2x}=(\Gamma_{0,0}+\Gamma_{3,0}+\Gamma_{0,1}-\Gamma_{3,1})/2$, ${\cal{P}}=\Gamma_{0,0}$, and ${\cal T}=\Gamma_{0,0}{\cal K}$ with $\Gamma_{i,j}=\sigma_{i}\otimes\sigma_{j}$ ($i,j=0,1,2,3,\pm$), Here, $\bm{\sigma}$ denote the Pauli matrixes, $\sigma_{0}$ is the identity matrix, and $\sigma_{\pm}=\sigma_{1}\pm i\sigma_{2}$. The corresponding model Hamiltonian under these symmetry constraints may be written as [59, 60]
${\cal H}_{0}$ $=\left[\begin{array}[]{cc}H_{1}&H^{\prime}\\ H^{\prime\dagger}&H_{2}\end{array}\right],$ (4)
where $H_{1}=t_{1}\cos\frac{k_{x}+k_{z}}{2}\sigma_{1}$, $H_{2}=t_{1}\cos\frac{k_{y}+k_{z}}{2}\sigma_{1}$ and
$H^{\prime}$ $=t_{2}\left[\begin{array}[]{cc}\cos\frac{k_{z}}{2}&\cos\frac{k_{y}}{2}\\ \cos\frac{k_{x}}{2}&\cos\frac{k_{x}+k_{y}+k_{z}}{2}\end{array}\right],$ (7)
with $t_{i}$ ($i=1,2$) the real parameters.
Then, we introduce the spin degree of freedom and turn on a AM ordering along $z$-direction [see Fig. 2(a)]. The AM ordering breaks ${\cal T}$, $\tilde{{\cal C}}_{4z}$ and $C_{2x}$, while holds $\tilde{{\cal C}}_{2z}$, ${\cal{P}}$ and the combined operation ~$\tilde{{\cal M}}_{z}={\cal{P}}\tilde{{\cal C}}_{2z}$, $\tilde{{\cal C}}_{4z}{\cal T}$ and $C_{2x}{\cal T}$. To the leading order, the AM Hamiltonian may be written as [61]
${\cal H}$ $=$ ${\cal H}_{0}s_{0}+J_{z}s_{z}\Gamma_{3,0},$ (8)
with $\bm{s}$ the Pauli matrixes acting on spin space and $J_{z}$ denoting the strength of the AM potential. When the AM potential ($J_{z}>0$) is much stronger than other hopping and energy terms, the eight bands are divided into two groups with each group including four bands. Since the two groups are separated in energy space by $2J_{z}$, the model (8) can be simplified as two effectively four-band model [62]. The basis of the upper subsystem is $(|A\uparrow\rangle,|B\uparrow\rangle,|C\downarrow\rangle,|D\downarrow\rangle)^{T}$ and that for the lower subsystem is $(|A\downarrow\rangle,|B\downarrow\rangle,|C\uparrow\rangle,|D\uparrow\rangle)^{T}$. We focus on the upper subsystem and omit $J_{z}$ from it, as it only shifts the bands. Then the upper four-band model can be expressed as [63]
${\cal H}_{\rm{eff}}$ $=\frac{t^{2}_{2}}{4J_{z}}+\left[\begin{array}[]{cc}H_{1}&0\\ 0&H_{2}\end{array}\right]+\frac{t^{2}_{2}}{4J_{z}}\left[\begin{array}[]{cc}H_{1}^{\prime}&0\\ 0&H_{2}^{\prime}\end{array}\right],$ (13)
with
$H_{1}^{\prime}({\bm{k}})$ $=2(\cos\frac{k_{x}}{2}\cos\frac{k_{z}}{2}+\cos\frac{k_{y}}{2}\cos\frac{k_{x}+k_{y}+k_{z}}{2})\sigma_{1}$
$+\left[\begin{array}[]{cc}\cos k_{y}+\cos k_{z}&0\\ 0&\cos k_{x}+\cos(k_{x}+k_{y}+k_{z})\end{array}\right],$ (16)
and $H_{2}^{\prime}(k_{x},k_{y},k_{z})=H_{1}^{\prime}(k_{y},k_{x},k_{z})$.
The band structure of ${\cal H}_{\rm{eff}}$ is calculated and plotted in Ref. [57], which qualitatively reproduces the band structure and the AM $Z^{3}$ nodal net of the $\beta$-Fe2(PO4)O, as well as the Q1D spin transport. Moreover, we find that the Q1D spin transport will disappear when the AM $Z^{3}$ nodal net is removed [57]. Our simple four-band effective model not only further demonstrate the closed relationship between AM $Z^{3}$ nodal net and Q1D spin transport, but also can serve as a good starting point for the subsequent study of the physics of AM $Z^{3}$ nodal net.
Conclusions.–We have demonstrated the feasibility of achieving ideal AM $Z^{3}$ nodal net in real materials and shown that the novel Q1D spin transport phenomena inherent to this topological structure. We have detailed the mechanisms whereby electrons with up-spin predominantly flow in one direction, while those with down-spin move primarily in a perpendicular direction, a characteristic distinctly contrasting with conventional 3D bulk materials and Q1D chain materials. This behavior could be effectively detected through the measurement of anisotropic spin resistance, offering a practical approach to exploring the unique physics of AM systems. Our work not only provides a solid foundation for exploring the novel physics of AM systems but also holds great promise for applications in topological spintronics.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (Grants Nos. 12004035, 12274112, 12347117 and 12304188) and the China Postdoctoral Science Foundation (Grants Nos. 2024M754079 and GZC20242181). X. Zhang also thanks the sponsorship from S&$\&T$ Program of Hebei (225676163GH). Appendix on numerical methods.– In this work, the electronic conductivity $\sigma(\mu,T)$ was calculated using the Boltzmann transport equation [64]. The conductivities for the two spin channels have a strong dependence on the chemical potential $\mu$ and temperature $T$, which can be calculated as follows [65, 66]:
[…]
where $f(E,\mu,T)$ is the usual Fermi-Dirac distribution function, defined as
$f(E,\mu,T)=\frac{1}{e^{(E-\mu)/k_{B}T}+1},$ (18)
and $\Sigma_{ij}(\epsilon)$ is the transport distribution function tensor, expressed as
$\Sigma_{ij}(\epsilon)=\frac{1}{N_{k}\Omega_{c}}\sum_{n,\mathbf{k}}v_{i}(n,\mathbf{k})v_{j}(n,\mathbf{k})\tau(n,\mathbf{k})\delta(E-E_{n,\mathbf{k}}),$ (19)
where $n$ is the band index, $E_{n,\mathbf{k}}$ is energy of the $n$-th band at momentum point $\mathbf{k}$, $v_{i}(n,\mathbf{k})$ denotes the $i$-th component of the velocity of the $n$-th band at $\mathbf{k}$, $\delta$ stands for Dirac’s delta function, and $\tau(n,\mathbf{k})$ is the relaxation time. $N_{k}$ and $\Omega_{c}$ represent the number of $k$-points used to sample the Brillouin zone and the unit cell volume, respectively. In this work, the relaxation time approximation was adopted [66, 67, 68], in such case $\tau$ is independent of $n$ and $\mathbf{k}$.
For $\beta$-Fe2(PO4)O, the electronic conductivity was calculated at a temperature of 50 K, which is much lower than Néel temperature of the system $T_{N}$ = 408 K, using a $100\times 100\times 100$ momentum mesh via the Wannier90 code [69, 70]. Since the relaxation time of the typical AM metal RuO2 is theoretically estimated as $\sim$ 600 fs [41, 71], we set the relaxation time of $\beta$-Fe2(PO4)O as a constant of 500 fs, which is comparable with that of RuO2. Similarly, for LiTi2O4 and Co2(PO4)O compounds, $T=50$ K, $\tau=500$ fs, and a momentum mesh of $100\times 100\times 100$ are used to calculate the electronic conductivity.